# DiffusionTrend: A Minimalist Approach to Virtual Fashion Try-On

### v1

@@ Abstract
Keywords Diffusion Model  $\cdot$ Virtual Try-On  $\cdot$ Image Editing
1 Introduction
Vitrual try-on technology Song et al. (2023); Islam et al. (2024), which digitally superimposes images of models wearing various outfits, represents a pivotal innovation in the fashion sector. This advancement offers consumers an immersive and interactive experience with garments, allowing them to preview how clothing might look on them without physically visiting stores. For retailers, virtual try-on technology streamlines their operations by obviating the need to employ live models for merchandise display. It also circumvents the financial burdens associated with traditional product photography, leading to a marked improvement in operational efficiency. For commercial platforms, virtual try-on technology serves as a magnet for attracting a larger user base and fostering user loyalty. Collectively, virtual try-on is reshaping the landscape of the apparel retail industry, propelling the fashion sector toward a future characterized by heightened efficiency and customization.
Table 1: Comparison of input requirements for previous virtual try-on models. A checkmark (✓) indicates that the input modality is required, while a dash (-) indicates that it is not or not mentioned.
@@ 1 Introduction
TryOnDiffusion Zhu et al. (2023) - -
DCI-VTON Gou et al. (2023) -
LaDI-VTON Morelli et al. (2023) - -
WarpDiffusion Li et al. (2023a) - - -
OOTDiffusion Xu et al. (2024) - -
StableVITON Kim et al. (2023) - -
IDM-VTON Choi et al. (2024) - -
Wear-Any-Way Chen et al. (2024) - -
@@
Traditional virtual try-on solutions Choi et al. (2021); Lee et al. (2022); Ge et al. (2021); Xie et al. (2023) are predicated on a two-stage pipeline utilizing Generative Adversarial Networks (GANs) Goodfellow et al. (2014). The initial stage in this framework involves the application of an explicit warp module to deform the clothing to the desired area on the body. The subsequent stage integrates the deformed clothing using a GAN-based try-on generator. To attain accurate clothing deformation, earlier methodologies Bai et al. (2022); Ge et al. (2021); Han et al. (2019); Lee et al. (2022); Xie et al. (2023) have employed a trainable network designed to estimate a dense flow map Zhou et al. (2016), thereby facilitating the mapping of the clothing onto the human form. In parallel, various approaches Choi et al. (2021); Ge et al. (2021); Xie et al. (2023); Lee et al. (2022); Yang et al. (2020); Issenhuth et al. (2020), have been proposed to address the misalignment issues between the warped clothing and the human body. Techniques such as normalization Choi et al. (2021) and distillation Ge et al. (2021); Issenhuth et al. (2020) have been implemented to mitigate these discrepancies. Recent advancements in diffusion models Ho et al. (2020) have led to a notable enhancement in the quality of image synthesis tasks, with a particular emphasis on the domain of virtual try-on. In this context, contemporary research endeavors have leveraged pre-trained text-to-image diffusion models to produce high-fidelity results. The TryOnDiffusion model Zhu et al. (2023), for instance, employs a dual U-Net architecture to perform the try-on task. LADI-VTON Morelli et al. (2023) and DCI-VTON Gou et al. (2023) either conceptualize clothing items as pseudo-words or integrate garments through the use of warping networks into pre-trained diffusion models.
@@
Despite these advancements, a review of existing studies reveals that the high precision and realism achieved by current methods require training on extensive try-on datasets Morelli et al. (2022), particularly in diffusion-based approaches Zhu et al. (2023); Morelli et al. (2023); Gou et al. (2023). Diffusion models simulate the data generation process by iteratively modeling the diffusion and reverse diffusion steps. Each iteration involves complex probability distribution calculations and requires substantial computational resources, leading to considerable overall training costs. In the words, current methods rely on resource-intensive network training. While intensive training is well-suited for generating complex poses and capturing fine image details, the substantial computational costs can make it an unattractive option. Besides, as shown in Table 1, these methods often require multiple auxiliary modules or APIs as model inputs, such as densepose Güler et al. (2018), segmentation maps Gong et al. (2017), clothes-agnostic images/masks/representations Han et al. (2018) and keypoints Cao et al. (2017). The integration of these tools not only increases inference overhead but also suffers from insufficient encapsulation in some cases, leaving non-professional users to deal with complex configurations and unstandardized interfaces, which significantly hampers accessibility and user experience. There is a pressing need for the community to explore more accessible and less resource-intensive methods, with the effort of less compromise on the quality of image generation—a domain that, to our knowledge, remains uncharted to date.
@@
In marked contrast, our proposed DiffusionTrend offers a streamlined, lightweight training approach that circumvents the need to retrain a diffusion model, thus liberating it from the reliance on expensive and resource-intensive computational infrastructure. This technique reduces inference overhead and simplifies the workflow by dispensing with the requirement for intricate segmentation, pose extraction, and other preliminary processing steps for the input images, leading to a more accessible and economical solution. Specifically, a lightweight, compact CNN is utilized to outline the clothing in both the model and garment images. It harmonizes image and textual features and conducts clustering at the feature level to produce effective masks. Subsequently, we make full use of the latents derived from DDIM inversion Song et al. (2020), which are replete with prior information and act as superior conduits for the detailed features of the garment. During the early stages of the diffusion denoising process, the seamless integration of the target garment into the model’s image reconstruction is achieved by blending the latent representations of both the model and the garment. In the subsequent stages of denoising, leveraging the self-repairing capabilities of the pre-trained latent diffusion model, we maintain the model’s identity and background coherence by selectively replacing the latents in the background areas. This approach ensures that the garment is harmoniously merged into the overall image. Fig. 1 shows several examples of our proposed DiffusionTrend.
@@
Figure 1: Images generated by the proposed DiffusionTrend model, given an input target model and a try-on clothing item both from DressCode dataset Morelli et al. (2022).
2 Related Work
2.1 Image Editing through Diffusion Processes
@@
The practice of integrating specific content into a base image to produce realistic composites is prevalent in image editing leveraging diffusion processes. Initially, the field was dominated by text-based models for image editing Brooks et al. (2023); Kawar et al. (2023). InstructPix2Pix Brooks et al. (2023) employs paired data to train diffusion models capable of generating an edited image from an input image and a textual instruction. Conversely, Imagic Kawar et al. (2023) harnesses a pre-trained text-to-image diffusion model to generate text embeddings that align with both the input image and the target text. The abstract nature of text poses a limitation in accurately delineating the subtleties of objects, therefore, image conditioning was introduced to offer more concrete and precise descriptions. DCFF Xue et al. (2022) pioneers the use of pyramid filters for image composition, which was subsequently advanced by Paint by Example Yang et al. (2023), employing CLIP embeddings of the reference image to condition the diffusion model. Most contemporary methodologies, such as Dreambooth Ruiz et al. (2023) (all model parameters), Textual Inversion Gal et al. (2022) (a word vector for novel concepts), and Custom-Diffusion Kumari et al. (2023) (cross-attention parameters), rely heavily on fine-tuning techniques. In contrast, a handful of approaches Hertz et al. (2022); Cao et al. (2023) adopts a training-free paradigm. Prompt-to-prompt Hertz et al. (2022) modifies the input text prompt to steer the cross-attention mechanism for nuanced image editing, while MasaCtrl Cao et al. (2023) employs a mask-guided mutual attention strategy for non-rigid image synthesis and editing. These training-free methods offer a cost-effective alternative, eliminating extensive training while still delivering commendable generative outcomes.
2.2 Virtual Try-on with Diffusion Models
@@
Diffusion models have demonstrated remarkable efficacy in the domain of image editing, with image-based virtual try-on representing a specialized subset of these tasks, contingent upon a specific garment image. Adapting text-to-image diffusion models to accommodate images as conditions, is straightforward, but the spatial discrepancies between the garment and the subject’s pose challenge the fidelity of texture details in the virtual try-on outcomes Li et al. (2023a); Gou et al. (2023); Morelli et al. (2023). Methodologies such as WarpDiffusion Li et al. (2023a), DCI-VTON Gou et al. (2023), and LADI-VTON Morelli et al. (2023) conceptualize clothing as pseudo-words, employing warping techniques through CNN networks to adjust clothing to various poses, yielding satisfactory results. TryOnDiffusion Zhu et al. (2023) employs a dual U-Net architecture for the virtual try-on task, implicitly conducting garment warping through the interplay between cross-attention layers. This approach effectively resolves the issue of texture misalignment without the need for a dedicated warp module. Similarly, the StableVITON Kim et al. (2023) incorporates zero cross-attention blocks to condition the intermediate feature maps of a spatial encoder, thereby circumventing the requirement for a warp module. The Wear-Any-Way Chen et al. (2024) enhances the process of virtual garment alteration, providing more adaptable control over the manner in which clothing is depicted. The IDM-VTON Choi et al. (2024) enhances the virtual try-on process by integrating attention mechanisms and high-level semantic encoding into the diffusion model.
3 Methodology
3.1 Preliminaries
@@
DDIM Inversion. DDIM inversion uses DDIM sampling Song et al. (2020) to ensure deterministic sampling by setting the variance in Eq. (2). It assumes the reversibility of the ordinary differential equation Chen et al. (2018) through incremental steps, ensuring a controlled transition from initial state $z_{0}$ to final noise $z_{T}$:
$z^{*}_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}z^{*}_{t-1}-\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}\Bigg{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\Bigg{)}\cdot\epsilon_{\theta}\big{(}z^{*}_{t-1},t-1,\tau_{\theta}(\mathcal{P})\big{)},$ (3)
@@
To address the challenge of extensive training overhead and various model inputs, we choose not to alter any weights or structures of the pre-trained diffusion model. Instead, we introduce a lite-training visual try-on method called “DiffusionTrend”. Our methodology consists of three stages. 1) A lightweight and compact CNN accurately delineates the apparel in both the model and garment images. 2) At an appropriate point in the process, garment details are integrated into the reconstruction phase of the model image. 3) To ensure the coherence of the generated background with the model, we use a latent substitution technique to restore the background, leveraging the diffusion model’s restorative properties to blend it seamlessly with the newly rendered apparel. A comprehensive discussion of the first stage is in Sec. 3.2, while the latter two stages are discussed in Sec. 3.3.
Figure 2: Inference cost comparison: Segment Anything Model (SAM) Kirillov et al. (2023) vs. our Apparel Localization Network.
3.2 Precise Apparel Localization
In conventional GAN-based Lee et al. (2022); Ge et al. (2021); Xie et al. (2023) or current Diffusion-based Zhu et al. (2023); Gou et al. (2023); Morelli et al. (2023) try-on methods, a precise clothing mask is crucial. This mask ensures correct apparel placement on the model and accurate extraction of garment features while avoiding interference from non-garment regions.
Traditional segmentation models He et al. (2017); Kirillov et al. (2023) excel in mask generation tasks. However, after thorough evaluation, we have chosen not to employ pre-built models like Segment Anything Model (SAM) Kirillov et al. (2023). Instead, we propose a lightweight CNN tailored for our application. This decision is informed by the following considerations:
@@ 3.2 Precise Apparel Localization
Task-Specific Optimization Over General-Purpose Segmentation: Off-the-shelf models like SAM are engineered for a wide range of semantic segmentation tasks and lack customization for the nuances of virtual try-on applications. Our proposed lightweight CNN, on the other hand, incorporates task-specific textual cues (e.g., “top", “skirt") to enable precise segmentation of clothing areas, sidestepping irrelevant regions.
@@
An apparel-related prompt $\mathcal{P}$ is transformed into a text embedding $\mathcal{T}_{0}=Clip(\mathcal{P})$ by a $Clip$ Radford et al. (2021) text encoder with fixed parameters. The text embedding $\mathcal{T}_{0}$ is then advanced through two $FC_{i}$ layers with a $ReLU$ function in between, to produce the text-derived features $\mathcal{T}_{2}$ as:
@@
Our apparel localization network is notably lightweight, making the training process highly cost-effective when compared to methods Li et al. (2023a); Kim et al. (2023); Gou et al. (2023) that require training a diffusion model. It takes us only 20 hours to process the entire DressCode dataset Morelli et al. (2022) using two RTX 3090 GPUs. As shown in Fig. 2, our proposed localization network operates fully automatically with only 2.00 GFLOPs and 0.15G of memory, and its accuracy analysis can be found in Sec. 4.3.
@@
3.3 Virtual Try-On Image Reconstruction
@@
Garment Latent Infusion. Recent studies Choi et al. (2024); Xu et al. (2024) indicate that employing U-Net for feature extraction necessitates treating it as a high-parameter module, which entails considerable training expenses. This approach conflicts with our research goal of achieving a training-free diffusion model solution, as such high computational costs undermine the feasibility of lightweight virtual try-on applications.
We first experiment with cross-attention-based MasaCtrl Cao et al. (2023), which involves a reference U-Net extracting garment features and performing key-value exchanges with the main U-Net during the attention stage. Although this method avoids additional training, it heavily relies on dual-branch U-Nets and attempts to guide attention interactions with masks derived from our apparel localization network. As shown in Fig. 3(b), the results fall short in accurately rendering fine-grained garment details. We surmise that direct feature injection through cross-attention, while effective at generating semantically coherent content, lacks the ability to capture the intricate details essential for high-quality virtual try-on results.
Figure 5: Comparison of results using explicit warping methods and perspective transformations. Left: garment with warped/perspective transformation; Right: results generated using the processed garment.
@@
Given a model image $\mathcal{I}^{m}$ and a garment image $\mathcal{I}^{g}$, we utilize our apparel localization network to acquire the masks $\mathcal{M}^{m}$ and $\mathcal{M}^{g}$. A natural thought is to leverage an explicit warp module to align the in-store garment $\mathcal{I}^{g}$ with the pose, position, and content of the clothing in $\mathcal{I}^{m}$. Initially, we adopted the warp module from DCI-VTON Gou et al. (2023) to preprocess garment images. However, the results were unsatisfactory. In Fig. 5, the warped garment images introduce severe distortions into the inversion and reconstruction phases. This, in turn, leads to highly distorted virtual try-on results, significantly degrading the overall quality of the pipeline.
We own this to incompatibility between the independent warp module and the untrained U-Net in our training-free paradigm. In the absence of collaborative training, the independent explicit warp module tends to produce incomplete warped images. Starting the inversion process with such incomplete warped images further hinders reconstruction quality, as the pipeline is unable to recover from these distortions.
Consequently, we opt for perspective transformations to address the alignment problem. By computing the bounding boxes $\mathcal{B}^{m}$ and $\mathcal{B}^{g}$ for $\mathcal{M}^{m}$ and $\mathcal{M}^{g}$, respectively, we reshape the in-store garment image $\mathcal{I}^{g}$ to align with the position and size of the clothing in the model image $\mathcal{I}^{m}$.
@@
Once aligned, the model image $\mathcal{I}^{m}$ and the adjusted garment image $\mathcal{I}^{g}$ are transformed into their latent representations $z_{0}^{m}=\mathcal{E}(\mathcal{I}^{m})$ and $z_{0}^{g}=\mathcal{E}(\mathcal{I}^{g})$, where $\mathcal{E}$ denotes the encoder. These latents are then subjected to the DDIM Inversion process in Eq. (3), yielding noisy latent sets $\{z^{*m}_{t}\}_{t=0}^{T}$ and $\{z^{*g}_{t}\}_{t=0}^{T}$.
Starting from $z^{*m}_{T}$, we reconstruct the model image using Eq. (2) where the latent at the $t$-th time step is $z_{t}^{m}$. We aim to decode an image of the model wearing the garment shown in image $\mathcal{I}^{g}$, achieving a harmonious fusion of the model’s appearance with the desired attire. A literature review Wu et al. (2023) has shown that the latent derived from the inversion process is rich in prior information, making it ideal for capturing the intricate features of the target garment. An effective strategy is to integrate the garment latent $z_{t_{1}}^{*g}$, preserved during the inversion stage, into the core of the garment mask $\mathcal{M}^{g}$ at time step $t_{1}$. This occurs early in the denoising process as:
@@
The denoising continues with the infused latent. It ensures the seamless information transferred from the garment image to the model image, and achieves a precise attire on the model.
While perspective transformations cannot simulate the complex deformations in real try-on scenarios, they offer a simple and lightweight solution for basic alignment, aligning with our goal of a training-free virtual try-on framework and lay the groundwork for future advancements in garment alignment.
@@
Table 2: Quantitative results on the VITON-HD and DressCode datasets.

### v2

@@ Abstract
Index Terms:
Diffusion Model, Virtual Try-On, Image Editing
I Introduction
Virtual try-on technology [1, 2], which digitally superimposes images of models wearing various outfits, represents a pivotal innovation in the fashion sector. This advancement offers consumers an immersive and interactive experience with garments, allowing them to preview how clothing might look on them without physically visiting stores. For retailers, virtual try-on technology streamlines their operations by obviating the need to employ live models for merchandise display. It also circumvents the financial burdens associated with traditional product photography, leading to a marked improvement in operational efficiency. For commercial platforms, virtual try-on technology serves as a magnet for attracting a larger user base and fostering user loyalty. Collectively, virtual try-on is reshaping the landscape of the apparel retail industry, propelling the fashion sector toward a future characterized by heightened efficiency and customization.
TABLE I: Comparison of input requirements for previous virtual try-on models. A checkmark (✓) indicates that the input modality is required, while a dash (-) indicates that it is not or not mentioned.
@@ I Introduction
TryOnDiffusion [3] - -
DCI-VTON [4] -
LaDI-VTON [5] - -
WarpDiffusion [6] - - -
OOTDiffusion [7] - -
StableVITON [8] - -
IDM-VTON [9] - -
Wear-Any-Way [10] - -
@@
Traditional virtual try-on solutions [11, 12, 13, 14, 15, 16] are predicated on a two-stage pipeline utilizing Generative Adversarial Networks (GANs) [17]. The initial stage in this framework involves the application of an explicit warp module [18, 19, 20, 21, 22] to deform the clothing to the desired area on the body. The subsequent stage integrates the deformed clothing using a GAN-based try-on generator. To attain accurate clothing deformation, earlier methodologies [23, 13, 24, 12, 14, 25, 26] have employed a trainable network designed to estimate a dense flow map [27, 28, 29], thereby facilitating the mapping of the clothing onto the human form. In parallel, various approaches [11, 13, 14, 12, 30, 31, 32, 33], have been proposed to address the misalignment issues between the warped clothing and the human body. Techniques such as normalization [11] and distillation [13, 31] have been implemented to mitigate these discrepancies. Recent advancements in diffusion models [34] have led to a notable enhancement in the quality of image synthesis tasks, with a particular emphasis on the domain of virtual try-on. In this context, contemporary research endeavors have leveraged pre-trained text-to-image diffusion models to produce high-fidelity results. The TryOnDiffusion model [3], for instance, employs a dual U-Net architecture to perform the try-on task. LADI-VTON [5] and DCI-VTON [4] either conceptualize clothing items as pseudo-words or integrate garments through the use of warping networks into pre-trained diffusion models.
Figure 1: Images generated by the proposed DiffusionTrend model, given an input target model and a try-on clothing item both from DressCode dataset [35].
@@
Despite these advancements, a review of existing studies reveals that the high precision and realism achieved by current methods require training on extensive try-on datasets [35], particularly in diffusion-based approaches [3, 5, 4]. Diffusion models simulate the data generation process by iteratively modeling the diffusion and reverse diffusion steps. Each iteration involves complex probability distribution calculations and requires substantial computational resources, leading to considerable overall training costs. In the words, current methods rely on resource-intensive network training. While intensive training is well-suited for generating complex poses and capturing fine image details, the substantial computational costs can make it an unattractive option. Besides, as shown in Table I, these methods often require multiple auxiliary modules or APIs as model inputs, such as densepose [36], segmentation maps [37], clothes-agnostic images/masks/representations [38] and keypoints [39]. The integration of these tools not only increases inference overhead but also suffers from insufficient encapsulation in some cases, leaving non-professional users to deal with complex configurations and unstandardized interfaces, which significantly hampers accessibility and user experience. There is a pressing need for the community to explore more accessible and less resource-intensive methods, with the effort of less compromise on the quality of image generation—a domain that, to our knowledge, remains uncharted to date.
@@
In marked contrast, our proposed DiffusionTrend offers a streamlined, lightweight training approach that circumvents the need to retrain a diffusion model, thus liberating it from the reliance on expensive and resource-intensive computational infrastructure. This technique reduces inference overhead and simplifies the workflow by dispensing with the requirement for intricate segmentation, pose extraction, and other preliminary processing steps for the input images, leading to a more accessible and economical solution. Specifically, a lightweight, compact CNN is utilized to outline the clothing in both the model and garment images. It harmonizes image and textual features and conducts clustering at the feature level to produce effective masks. Subsequently, we make full use of the latents derived from DDIM inversion [40], which are replete with prior information and act as superior conduits for the detailed features of the garment. During the early stages of the diffusion denoising process, the seamless integration of the target garment into the model’s image reconstruction is achieved by blending the latent representations of both the model and the garment. In the subsequent stages of denoising, leveraging the self-repairing capabilities of the pre-trained latent diffusion model, we maintain the model’s identity and background coherence by selectively replacing the latents in the background areas. This approach ensures that the garment is harmoniously merged into the overall image. To further enhance robustness under diverse generation conditions, we apply an adaptive sampling termination strategy based on perceptual color difference to automatically determine the optimal number of sampling steps during inference. Fig. 1 shows several examples of our proposed DiffusionTrend.
@@
II Related Work
II-A Image Editing through Diffusion Processes
@@
The practice of integrating specific content into a base image to produce realistic composites is prevalent in image editing leveraging diffusion processes. Initially, the field was dominated by text-based models for image editing [41, 42]. InstructPix2Pix [41] employs paired data to train diffusion models capable of generating an edited image from an input image and a textual instruction. Conversely, Imagic [42] harnesses a pre-trained text-to-image diffusion model to generate text embeddings that align with both the input image and the target text. The abstract nature of text poses a limitation in accurately delineating the subtleties of objects, therefore, image conditioning was introduced to offer more concrete and precise descriptions. DCFF [43] pioneers the use of pyramid filters for image composition, which was subsequently advanced by Paint by Example [44], employing CLIP embeddings of the reference image to condition the diffusion model. Most contemporary methodologies, such as Dreambooth [45] (all model parameters), Textual Inversion [46] (a word vector for novel concepts), and Custom-Diffusion [47] (cross-attention parameters), rely heavily on fine-tuning techniques. In contrast, a handful of approaches [48, 49] adopts a training-free paradigm. Prompt-to-prompt [48] modifies the input text prompt to steer the cross-attention mechanism for nuanced image editing, while MasaCtrl [49] employs a mask-guided mutual attention strategy for non-rigid image synthesis and editing. These training-free methods offer a cost-effective alternative, eliminating extensive training while still delivering commendable generative outcomes.
II-B Virtual Try-on with Diffusion Models
@@
Diffusion models have demonstrated remarkable efficacy in the domain of image editing, with image-based virtual try-on representing a specialized subset of these tasks, contingent upon a specific garment image. Adapting text-to-image diffusion models to accommodate images as conditions, is straightforward, but the spatial discrepancies between the garment and the subject’s pose challenge the fidelity of texture details in the virtual try-on outcomes [6, 4, 5]. Methodologies such as WarpDiffusion [6], DCI-VTON [4], and LADI-VTON [5] conceptualize clothing as pseudo-words, employing warping techniques through CNN networks to adjust clothing to various poses, yielding satisfactory results. TryOnDiffusion [3] employs a dual U-Net architecture for the virtual try-on task, implicitly conducting garment warping through the interplay between cross-attention layers. This approach effectively resolves the issue of texture misalignment without the need for a dedicated warp module. Similarly, the StableVITON [8] incorporates zero cross-attention blocks to condition the intermediate feature maps of a spatial encoder, thereby circumventing the requirement for a warp module. The Wear-Any-Way [10] enhances the process of virtual garment alteration, providing more adaptable control over the manner in which clothing is depicted. The IDM-VTON [9] enhances the virtual try-on process by integrating attention mechanisms and high-level semantic encoding into the diffusion model.
III Methodology
III-A Preliminaries
@@
DDIM Inversion. DDIM inversion uses DDIM sampling [40] to ensure deterministic sampling by setting the variance in Eq. (2). It assumes the reversibility of the ordinary differential equation [50] through incremental steps, ensuring a controlled transition from initial state $z_{0}$ to final noise $z_{T}$:
$z^{*}_{t}=$ $\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}z^{*}_{t-1}-\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}\Bigg{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\Bigg{)}$
$\cdot\epsilon_{\theta}\big{(}z^{*}_{t-1},t-1,\tau_{\theta}(\mathcal{P})\big{)},$ (3)
@@
To address the challenge of extensive training overhead and various model inputs, we choose not to alter any weights or structures of the pre-trained diffusion model. Instead, we introduce a lite-training visual try-on method called “DiffusionTrend”. Our methodology consists of three stages. 1) A lightweight and compact CNN accurately delineates the apparel in both the model and garment images. 2) At an appropriate point in the process, garment details are integrated into the reconstruction phase of the model image. 3) To ensure the coherence of the generated background with the model, we use a latent substitution technique to restore the background, leveraging the diffusion model’s restorative properties to blend it seamlessly with the newly rendered apparel. A comprehensive discussion of the first stage is in Sec. III-B, while the latter two stages are discussed in Sec. III-C.
III-B Precise Apparel Localization
In conventional GAN-based [12, 13, 14] or current Diffusion-based [3, 4, 5] try-on methods, a precise clothing mask is crucial. This mask ensures correct apparel placement on the model and accurate extraction of garment features while avoiding interference from non-garment regions.
Figure 2: Inference cost comparison: Segment Anything Model (SAM) [51] vs. our Apparel Localization Network.
@@ III-B Precise Apparel Localization
Traditional segmentation models [52, 51] excel in mask generation tasks. However, after thorough evaluation, we have chosen not to employ pre-built models like Segment Anything Model (SAM) [51]. Instead, we propose a lightweight CNN tailored for our application. This decision is informed by the following considerations:
@@
Task-Specific Optimization Over General-Purpose Segmentation: Off-the-shelf models like SAM are engineered for a wide range of semantic segmentation tasks and lack customization for the nuances of virtual try-on applications. Our proposed lightweight CNN, on the other hand, incorporates task-specific textual cues (e.g., “top”, “skirt”) to enable precise segmentation of clothing areas, sidestepping irrelevant regions. We do not adopt the CLIP [53] image encoder, as it is primarily optimized for global image-text alignment and lacks the spatial resolution required for pixel-accurate apparel segmentation. Moreover, it introduces additional inference cost, which contradicts our lightweight design principle.
@@
An apparel-related prompt $\mathcal{P}$ is transformed into a text embedding $\mathcal{T}_{0}=Clip(\mathcal{P})$ by a $Clip$ [53] text encoder with fixed parameters. The text embedding $\mathcal{T}_{0}$ is then advanced through two $FC_{i}$ layers with a $ReLU$ function in between, to produce the text-derived features $\mathcal{T}_{2}$ as:
@@
To further enhance the precision and boundary regularity of the K-means-generated masks, we apply a classic active contour-based [54] refinement method as a post-processing step. This technique iteratively optimizes the mask boundary by minimizing an energy function that balances edge attraction and contour smoothness, producing more natural and connected garment regions. A comparative study of garment mask refinement strategies is presented in Sec. IV-C.
Our apparel localization network is notably lightweight, making the training process highly cost-effective when compared to methods [6, 8, 4] that require training a diffusion model. It takes us only 20 hours to process the entire DressCode dataset [35] using two RTX 3090 GPUs. As shown in Fig. 2, our proposed localization network operates fully automatically with only 2.00 GFLOPs and 0.15G of memory, and its accuracy analysis can be found in Sec. IV-C.
@@
III-C Virtual Try-On Image Reconstruction
@@
Garment Latent Infusion. Recent studies [9, 7] indicate that employing U-Net for feature extraction necessitates treating it as a high-parameter module, which entails considerable training expenses. This approach conflicts with our research goal of achieving a training-free diffusion model solution, as such high computational costs undermine the feasibility of lightweight virtual try-on applications.
We first experiment with cross-attention-based MasaCtrl [49], which involves a reference U-Net extracting garment features and performing key-value exchanges with the main U-Net during the attention stage. Although this method avoids additional training, it heavily relies on dual-branch U-Nets and attempts to guide attention interactions with masks derived from our apparel localization network. As shown in Fig. 3(b), the results fall short in accurately rendering fine-grained garment details. We surmise that direct feature injection through cross-attention, while effective at generating semantically coherent content, lacks the ability to capture the intricate details essential for high-quality virtual try-on results.
@@
Given a model image $\mathcal{I}^{m}$ and a garment image $\mathcal{I}^{g}$, we utilize our apparel localization network to acquire the masks $\mathcal{M}^{m}$ and $\mathcal{M}^{g}$. A natural thought is to leverage an explicit warp module to align the in-store garment $\mathcal{I}^{g}$ with the pose, position, and content of the clothing in $\mathcal{I}^{m}$. Initially, we adopted three representative methods, including the warping modules from DCI-VTON [4], SCW-VTON [55], and PL-VTON [56], to preprocess garment images. However, the results were unsatisfactory. In Fig. 10, the warped garment images introduce severe distortions into the inversion and reconstruction phases. This, in turn, leads to highly distorted virtual try-on results, significantly degrading the overall quality of the pipeline.
We attribute this to incompatibility between the independent warp module and the untrained U-Net in our training-free paradigm, with detailed analysis provided in Sec. IV-C. Consequently, we opt for perspective transformations to address the alignment problem. By computing the bounding boxes $\mathcal{B}^{m}$ and $\mathcal{B}^{g}$ for $\mathcal{M}^{m}$ and $\mathcal{M}^{g}$, respectively, we reshape the in-store garment image $\mathcal{I}^{g}$ to align with the position and size of the clothing in the model image $\mathcal{I}^{m}$.
@@
Once aligned, the model image $\mathcal{I}^{m}$ and the adjusted garment image $\mathcal{I}^{g}$ are transformed into their latent representations $z_{0}^{m}=\mathcal{E}(\mathcal{I}^{m})$ and $z_{0}^{g}=\mathcal{E}(\mathcal{I}^{g})$, where $\mathcal{E}$ denotes the encoder. These latents are then subjected to the DDIM Inversion process in Eq. (III-A), yielding noisy latent sets $\{z^{*m}_{t}\}_{t=0}^{T}$ and $\{z^{*g}_{t}\}_{t=0}^{T}$.
Starting from $z^{*m}_{T}$, we reconstruct the model image using Eq. (2) where the latent at the $t$-th time step is $z_{t}^{m}$. Our goal is to decode an image of the model wearing the garment from image $\mathcal{I}^{g}$, thereby achieving a harmonious fusion of the model’s appearance and the desired attire. A literature review [57] has shown that latents derived from the inversion process retains rich prior information. This property makes it ideal for capturing the intricate features of the target garment. An effective strategy is to integrate the garment latent $z_{t_{1}}^{*g}$, preserved during the inversion stage, into the masked region $M^{g}$ at time step $t_{1}$. This occurs early in the denoising process as:
@@
The denoising continues with the infused latent. It ensures the seamless information transferred from the garment image to the model image, and achieves a precise attire on the model. We choose latent-space fusion over image-space masking, as it enables progressive integration of garment details and avoids visual artifacts caused by hard region cuts.
While perspective transformations cannot simulate the complex deformations in real try-on scenarios, they offer a simple and lightweight solution for basic alignment, which aligns with our goal of a training-free virtual try-on framework and lays the groundwork for future advancements in garment alignment.
@@
TABLE II: Quantitative results on the VITON-HD and DressCode datasets.